write an equation to show what happend when, to a buffer solution containing equimolar amounts of PO43- and HPO42-, we add.
a) H3O+
_____ + _____ ----> _____ + _____

b) OH-
_____ + _____ ----> _____+ _____
please do the math
Answer #1

The reactions that occur are written:

a) PO4-3 + H3O + = HPO4-2 + H2O

b) HPO4-2 + OH- = PO4-3 + H2O

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.